Shift summary for OWL shift 10/Jan/2000

Subject: Shift summary for OWL shift 10/Jan/2000
From: Vivian O'Dell (
Date: Fri Feb 04 2000 - 15:13:18 CST

                     Shift summary for owl shift 10/Jan/2000

On Shift: RK, AG

Hours of beam available at all intensities: 5.3
Hours of beam available at nominal intensity: 5.3
Hours of E799 data (special runs not included): 4.7
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

o Fresh start after L3 accept fraction got screwed up.

o Tevatron quench.


Tapes brought to FCC this shift: none.


Accesses during shift:
   RK and BH went to install UCAL scintillators (after tev quench)


Runs ended this shift:

15451 - 11.5M - BEAM_799
15452 - 5.8M - BEAM_799

Overview of shift:

Smooth E799 data-taking until the quench at 05:25, when we start paging
people from the list.

After tev quence, we could not do a drt_fresh_start.
We re-booted ktev3 and then all was well.


Description of Shift:

run spill time comment

15451 249 0210 L3 accept fraction is screwed up. Stop the run and
                        dart_fresh_start, but unsuccessfully. Logout/login
                        couple of times, and now it works.

15452 1 0242 Run 15453 is started.

15452 0525 Quench at D38U. Est. for beam 0700.
                        -> Stop the run, start paging people.

Items to Follow up on Next Shift

  Don't forget to hit "reset quartz" button after each KFREND.

This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:13:22 CST